OpenStudy (mimicutiepie1):

What is the value of the expression? (14 – 6) • (2 + 4) A. 6 B. 14 C. 4 D. 48 Help me .-. please...im tired I need some homework help ;c

OpenStudy (benji):

What do you think the answer is? If you don't have one, how far did you get?

jhonyy9 (jhonyy9):

Mimi step by step - ok. ? 14-6 = ?

OpenStudy (mimicutiepie1):

ok well 14-6=6

jhonyy9 (jhonyy9):

no look please again

OpenStudy (mimicutiepie1):

eh but isn't it subtraction...?

OpenStudy (benji):

Yes it is, 12-6=6. What is 14-6=?

OpenStudy (mimicutiepie1):

sorry its 8!

jhonyy9 (jhonyy9):

yes 8 and 2+4= ?

OpenStudy (mimicutiepie1):

6

OpenStudy (mathmale):

Order of Operations rules apply here. You MUST do the work shown within parentheses first. Would you do that now and share what you 've done? Not just the answer, but the actual work.

jhonyy9 (jhonyy9):

ok nice so now multiplie the 8 by 6 = ?

OpenStudy (mimicutiepie1):

48 my answer is 48 ?

OpenStudy (benji):

Nice

jhonyy9 (jhonyy9):

yes you are right with 48 good work

OpenStudy (mathmale):

would you please share your work? Answers without work are next to meaningless.

OpenStudy (mimicutiepie1):

yay

OpenStudy (mathmale):

Good for you. Now, why not go back and review Order of Operations Rules, so that next time you'll know how to solve problems like this one yourself.